Without graphing, determine whether the following pairs of lines are (a) parallel, (b) perpendicular, or (c) neither parallel nor perpendicular:
5x-6y=19
6x+5y=-30

Answer by ewatrrr(24785)


Hi
Decide whether the pair of lines is parallel, perpendicular or neither.
5x-6y=19 OR y = (5/6)x - 19/6 | m = 5/6
6x+5y=-30 OR y = (-6/5)x -6 | m = -6/5
Lines are perpendicular as 5/6*-6/5 = -1